OpenStudy (anonymous):

Find X. 3^x•5^x+1=45

OpenStudy (anonymous):

Subtract 1, you get 3^x (5^x) = 44 Now take logs of both sides, you get log[3^x (5^x)] = log 44, using law of logs for left side, you get log 3^x + log 5^x = log 44 Use another law of logs for left side, you get x(log 3) + x(log 5) = log 44 Now, factor out x on left side, and then divide.